Five hundred children participated in a field demonstration. Their heights averaged is 110 cm with a standard deviation of 6 cm. What is the probability that the height of a child picked at random is less than 104cm?


1
Expert's answer
2022-03-16T05:06:08-0400

"\mu=110,\sigma=6"

"P(X<104)=P(Z<\frac{x-\mu}{\sigma})"

"=P(Z<\frac{104-110}{6})=P(Z<-1)"

"=0.1587"